# On Safety and Liveness Filtering Using Hamilton-Jacobi Reachability Analysis

### v1

@@ Abstract
Hamilton-Jacobi (HJ) reachability-based filtering provides a powerful framework to co-optimize performance and safety (or liveness) for autonomous systems. Under this filtering scheme, a nominal controller is minimally modified to ensure system safety or liveness. However, the resulting controllers can exhibit abrupt switching and bang-bang behavior, which is not suitable for applications of autonomous systems in the real world. This work presents a novel, unifying framework to design safety and liveness filters through reachability analysis. We explicitly characterize the maximal set of control inputs that ensures safety (or liveness) at a given state. Different safety filters can then be constructed using different subsets of this maximal set along with a projection operator to modify the nominal controller. We use the proposed framework to design three safety filters, each balancing performance, computation time, and smoothness differently. The proposed filters can easily handle dynamics uncertainties, disturbances, and bounded control inputs. We highlight their relative strengths and limitations by applying these filters to autonomous navigation and rocket landing scenarios and on a physical robot testbed. We also discuss practical aspects associated with implementing these filters on real-world autonomous systems. Our research advances the understanding and potential application of reachability-based controllers on real-world autonomous systems.
@@ I Introduction
An alternative approach is safety filtering, where a nominal control input (often performance-oriented) is minimally adjusted (i.e., “filtered”) to satisfy the safety constraint. If the input is safe, it is used directly; otherwise, it is projected to a set of safe control inputs. This approach guarantees safety while optimizing performance whenever the system safety is not at risk. Akin to safety filters, one can also design liveness filters, which ensure that the system reaches a desired set of states (e.g., the landing pad for a rocket) within a set timeframe while optimizing the performance (e.g., minimizing the control energy).
@@
Hamilton-Jacobi (HJ) reachability analysis [17, 19] is a popular mechanism to design these filters due to its ability to ensure safety and liveness for dynamical systems with general nonlinear dynamics, control bounds, and disturbances [18, 4, 5]. In reachability analysis, one is interested in computing the backward reachable tube (BRT) of a dynamical system, i.e. the set of all initial states from which the system will eventually reach a target set, despite the worst case disturbance. Thus, if the target set represents the set of desirable states (e.g., goal states for a vehicle), the BRT represents the set of states from which liveness can be guaranteed. Conversely, if the target set represents the set of undesirable states (e.g., obstacles), the BRT contains states which are potentially unsafe for the system and should therefore be avoided. Alongside the BRT, reachability analysis provides a safety controller (respectively liveness controller) that will provably steer the system away from the target set (respectively towards the target set). A simple safety filter uses the BRT and safety controller: the system employs the nominal controller outside the BRT and switches to the reachability safety controller at the BRT boundary, ensuring constant safety. However, controllers from HJ reachability are often extremum seeking, resulting in a bang-bang behavior. This behavior along with a sudden switching between the nominal and reachability controllers leads to jittery state and control trajectories, which is often undesirable for real-world autonomous systems [24].
@@
An alternative approach to filtering involves using Control Barrier Functions (CBF) for safety and Control Lyapunov Functions (CLF) for liveness filtering. Both utilize Lyapunov-like conditions to ensure the forward invariance of a set [2], which then aids in creating a quadratic program (QP) for a smooth blending of nominal and safety (or liveness) controllers. However, despite recent progress [9, 22, 21, 26], constructing a valid Lyapunov or barrier function for general nonlinear systems with control bounds and disturbances remains a challenge. Some recent studies have addressed this by framing CBF synthesis as an HJ reachability problem, capitalizing on the constructive attributes of reachability methods for obtaining a valid CBF [9]. The resultant CBF can be used for smooth, QP-based controller synthesis. However, a few questions remain: can we derive smooth control laws directly from reachability analysis without constructing a CBF first? How do these safety filters compare in performance and computation? And can we design such control laws for liveness filtering?
@@
In this work, we answer the aforementioned questions in affirmative and introduce a general framework to design safety and liveness filters using reachability analysis. Under the proposed framework, any safety and liveness filter can be thought of as a combination of two components: (a) a projection set – a set of safe or live control inputs that the nominal controller will be projected to; and (b) a projection operator to align a nominal control input with this set. Drawing on this insight, our key idea is to use HJ reachability analysis to explicitly characterize the set of all possible control inputs at a state that guarantees system safety or liveness. Various safety filters can then be obtained by using different subsets of this maximal safe control set as the projection set, along with a projection operator. Building upon this framework, we propose three different safety and liveness filters based on reachability analysis: a least restrictive safety filter, a smooth least restrictive filter, and a smooth blending filter. The proposed filters are versatile, applicable to both liveness and safety, accommodating control bounds, adversarial disturbances, model uncertainties, and time-specific liveness and safety properties (e.g., reach the target set within 5 seconds). We further compare these filters on various metrics including performance, computation time, controller tuning, and control profile smoothness, and highlight the key trade-offs. Beyond these specific filters, our approach provides a unifying framework for existing reachability-based filters and designing new ones using diverse subsets of the maximal safe set and projection operators. We showcase our approach in two different applications inspired by rocket landing within a given landing window and safe autonomous blimp navigation, as well as on a physical robot testbed involving a wheeled robot navigating through a cluttered environment.
@@ II Problem Statement
Consider an autonomous system with state $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$ that evolves according to dynamics
$\dot{x}=f(x,u,d),$ (1)
where $u\in\mathcal{U}$ and $d\in\mathcal{D}$ are the control and disturbance of the system, respectively. $d$ can represent potential model uncertainties or an actual, adversarial exogenous input to the system. We assume the dynamics is uniformly continuous, bounded, and Lipschitz continuous in $x$ for fixed $u$ and $d$; such an assumption guarantees that a unique system trajectory would solve (1) [10] for all initial states. Finally, let $\xi_{x,t}^{\mathbf{u},\mathbf{d}}(\tau)$ denote the system state at time $\tau$, starting from the state $x$ at time $t$ under control input $\mathbf{u}(\cdot)$ and disturbance signal $\mathbf{d}(\cdot)$ while following the dynamics (1). A control input $\mathbf{u}(\cdot)$ is defined as a measurable function mapping from the time horizon to the set of admissible controls $\mathcal{U}$, and a disturbance signal is similarly defined.
[…]
@@
$f_{\text{rocket}}(x,u)=\dfrac{d}{dt}\begin{bmatrix}y\\ z\\ \theta\\ \dot{y}\\ \dot{z}\\ \dot{\theta}\end{bmatrix}=\begin{bmatrix}\dot{y}\\ \dot{z}\\ \dot{\theta}\\ \cos(\theta)u_{1}-\sin(\theta)u_{2}\\ \sin(\theta)u_{1}+\cos(\theta)u_{2}-g\\ \alpha u_{1}\end{bmatrix}$ (2)
@@
Here $y$, $z$ denotes the horizontal and vertical positions of the center of mass, $\theta$ denotes the rocket’s heading, and $g$ is the acceleration due to the gravity of Earth. The task of the liveness controller is to generate thrust inputs $u=[|u_{y}|<250,|u_{z}|<250]$ such that the system is able to reach the landing pad (target set) =$\mathcal{L}=\{(y,z):|y|<20,0<z<20\}$, within a time horizon of $T=1s$. For this particular example, we assume there are no external disturbances acting on the system.
@@
We first synthesize a nominal controller that attempts to take the system to $\mathcal{L}$. The nominal controller is a sampling-based model predictive controller (MPC), whose objective is to minimize the distance between center of mass position and the boundary of $\mathcal{L}$, and the control energy over the entire trajectory. Intuitively, the nominal controller can be considered as a performance control that tends to drive the system state to $\mathcal{L}$ while optimizing the amount of control effort. The MPC controller is implemented in a receding horizon fashion with a horizon of 1000 steps. We use a first-order Euler discretization of the continuous dynamics in (2) with timestep, $\delta$ = 0.0005s. The system trajectory and the control inputs under the nominal controller are shown in Fig. 3 in grey dashed line. Evidently, the nominal controller is insufficient in assuring the liveness of our system. Such a guarantee will require a more careful and timeconsuming process of designing the nominal controller. In this work, we will use a liveness filter that can refine this nominal controller to produce a controller with the desired liveness properties.
Running Example (Safety). To demonstrate effectiveness of the safety filters, we use a 4D system described by the following dynamics:
$\dot{x}=v\cos(\theta)+{d}_{1}$ (3)
$\dot{y}=v\sin(\theta)+{d}_{2}$
$\dot{z}=u_{1}$
$\dot{\theta}=u_{2},$
The state consists of the $X,Y$ and $Z$ positions and the heading on the $XY$ plane of the system. The control inputs are the vertical velocity in the $Z$ direction and the rate of change of heading. In addition, there are velocity disturbances in the $X$ and $Y$ directions.
We consider a $10m$ x $10m$ x $5m$ position state space (Fig. 2) where the objective is to reach a target area of radius $R=0.5m$ located at $(x,y,z)=(7,7,3)m$ (pink sphere) while avoiding collision with static obstacles represented as spheres of various sizes spread across the position state space. In (3) $v=2m/s$ is the velocity in the $XY$ plane, the control and disturbance bounds are given as $u_{1}\in[-1,1]m/s$, $u_{2}\in[-\pi,\pi]\text{rad}/s$ and $||\begin{bmatrix}d_{1}&d_{2}\end{bmatrix}^{\top}||_{2}\leq 0.5m/s$.
For this system we use an MPC-based nominal controller, whose objective is to minimize distance between the system and the target area, subjected to the system dynamics and the control bounds. To highlight the impacts of safety filtering, the nominal controller is built such that it does not consider obstacle avoidance in its objective. The system state trajectories and control profiles under the nominal controller are shown in grey in Fig. 7 and Fig. 8 respectively. As expected, the nominal controller fails to satisfy the safety constraint (i.e. avoiding the obstacles).
@@ III-A Liveness and Safety Problems in HJ Reachability
In HJ reachability analysis, the liveness and safety problems are posed as optimal control problems; informally, both problems intend to find a control signal $\mathbf{u}(\cdot)$ that steers the system as “deep” into or as “far away” from the target set as possible. To capture this semantic, the minimum distance between the system and the target set $\mathcal{L}$ over the time horizon is defined as the objective of the optimal control problems. Let $l:\mathcal{X}\rightarrow\mathbb{R}$ be some bounded and Lipschitz continuous function whose sub-zero level is given by the target set: =$\mathcal{L}=\{x:l(x)\leq 0\}$. Here, we present the reachability analysis for the liveness case and then comment on the safety case. Given $l$, the liveness problem is defined in (4):
$\begin{split}\max_{\mathbf{d}(\cdot)}\min_{\mathbf{u}(\cdot)}&\min_{\tau\in[t,T]}l(\xi_{x,t}^{\mathbf{u},\mathbf{d}}(\tau))\\ s.t.&\quad\dot{x}=f(x,u,d)\end{split}$ (4)
Where the minimum cost over time for the trajectory is defined as the cost function $J(x,t,\mathbf{u}(\cdot),\mathbf{d}(\cdot))=\min_{\tau\in[t,T]}l(\xi_{x,t}^{\mathbf{u},\mathbf{d}}(\tau))$ The above optimization problem (4) finds the minimum distance to the target set over the system trajectory, under the optimal control and the worst case disturbance. Thus, the system can reach the target set if and only if the minimum distance is less than 0, and this minimum distance is captured by the value function:
$V(x,t)=\max_{\mathbf{d}(\cdot)}\min_{\mathbf{u}(\cdot)}J(x,t,\mathbf{u}(\cdot),\mathbf{d}(\cdot))$ (5)
The value function (5) can be computed using dynamic programming, and it satisfies the following final value Hamilton-Jacobi-Issacs Variational Inequality (HJI-VI) [19, 16]:
[…]
$V(x,T)=l(x),\quad\text{for}\ t\in\left[0,T\right]$
Here, $D_{t}V(x,t)$ and $\nabla V(x,t)$ denote the temporal derivative and the spatial gradients of the value function $V(x,t)$, respectively. The Hamiltonian encodes how the control and disturbance interact with the system dynamics and is given as:
$H(x,t,\nabla V(x,t))=\min_{u\in\mathcal{U}}\max_{d\in\mathcal{D}}\nabla V(x,t)\cdot f(x,u,d)$ (7)
@@
$\mathcal{G}_{\text{live}}(t)\coloneqq\{x:V(x,t)\leq 0\}$ (8)

### v2

@@ Abstract
Hamilton-Jacobi (HJ) reachability-based filtering provides a powerful framework to co-optimize performance and safety (or liveness) for autonomous systems. Under this filtering scheme, a nominal controller is minimally modified to ensure system safety or liveness. However, the resulting controllers can exhibit abrupt switching and bang-bang behavior, which is not suitable for applications of autonomous systems in the real world. This work presents a novel, unifying framework to design safety and liveness filters through reachability analysis. We explicitly characterize the maximal set of control inputs that ensures safety (or liveness) at a given state. Different safety filters can then be constructed using different subsets of this maximal set along with a projection operator to modify the nominal controller. We use the proposed framework to design three safety filters, each balancing performance, computation time, and smoothness differently. We highlight their relative strengths and limitations by applying these filters to autonomous navigation and rocket landing scenarios and on a physical robot testbed. We also discuss practical aspects associated with implementing these filters on real-world autonomous systems. Our research advances the understanding and potential application of reachability-based controllers on real-world autonomous systems.
@@ I Introduction
An alternative approach is safety filtering, where a nominal control (often performance-oriented) is minimally adjusted (i.e., “filtered”) to satisfy the safety constraint. If the input is safe, it is used directly; otherwise, it is projected to a set of safe control inputs. This guarantees safety while optimizing performance whenever the system safety is not at risk. Akin to safety filters, one can also design liveness filters, which ensure that the system reaches a desired set of states (e.g., the landing pad for a rocket) within a set timeframe while optimizing the performance (e.g., minimizing the control energy).
@@
Hamilton-Jacobi (HJ) reachability analysis [18, 20] is a popular mechanism to design these filters due to its ability to ensure safety and liveness for dynamical systems with general nonlinear dynamics, control bounds, and disturbances [19, 3, 4]. In reachability analysis, one is interested in computing the backward reachable tube (BRT) of a dynamical system, i.e. the set of all initial states from which the system will eventually reach a target set, despite the worst case disturbance. Thus, if the target set represents the set of desirable states, the BRT represents the set of states from which liveness can be guaranteed. Conversely, if the target set represents the set of undesirable states, the BRT contains states which are potentially unsafe for the system and should be avoided. Alongside the BRT, reachability analysis provides a safety controller (respectively liveness controller) that will provably steer the system away from the target set (respectively towards the target set). A simple safety filter uses the BRT and safety controller: the system employs the nominal controller outside the BRT and switches to the reachability safety controller at the BRT boundary, ensuring constant safety. However, controllers from HJ reachability are often extremum seeking, resulting in a bang-bang behavior. This behavior along with a sudden switching between the nominal and reachability controllers leads to jittery state and control trajectories, which is often undesirable for real-world autonomous systems [25].
@@
An alternative approach to filtering uses Control Barrier Functions (CBF) for safety and Control Lyapunov Functions (CLF) for liveness filtering. Both utilize Lyapunov-like conditions to ensure the forward invariance of a set [1], which then aids in creating a quadratic program (QP) for a smooth blending of nominal and safety (or liveness) controllers. However, despite recent progress [10, 23, 22, 28], constructing a valid Lyapunov or barrier function for general nonlinear systems with control bounds and disturbances remains a challenge. Some recent studies have addressed this by framing CBF synthesis as an HJ reachability problem, capitalizing on the constructive attributes of reachability methods for obtaining a valid CBF [10]. The resultant CBF can be used for smooth, QP-based controller synthesis. However, a few questions remain: can we derive smooth control laws directly from reachability analysis without constructing a CBF first? How do these safety filters compare in performance and computation? And can we design such control laws for liveness filtering?
@@
In this work, we answer the aforementioned questions in affirmative and introduce a general framework to design safety and liveness filters using reachability analysis. Under the proposed framework, any safety and liveness filter can be thought of as a combination of two components: (a) a projection set – a set of safe or live control inputs that the nominal controller will be projected to; and (b) a projection operator to align a nominal control input with this set. Drawing on this insight, our key idea is to use HJ reachability analysis to explicitly characterize the set of all possible control inputs at a state that guarantees system safety or liveness. Various safety filters can then be obtained by using different subsets of this maximal safe control set as the projection set, along with a projection operator, to mathematically characterize the proposed framework we will use the notation presented in Table I. Building upon this approach, we propose three different safety and liveness filters based on reachability analysis: a least restrictive safety filter, a smooth least restrictive filter, and a smooth blending filter. The proposed filters are versatile, applicable to both liveness and safety, accommodating control bounds, adversarial disturbances, model uncertainties, and time-specific liveness and safety properties (e.g., reach the target set within 5 seconds). We further compare these filters on various metrics including performance, computation time, controller tuning, and control profile smoothness, as well as a comparison to CBF methods for the safety filter. Beyond these specific filters, our approach provides a unifying framework for existing reachability-based filters and designing new ones using diverse subsets of the maximal safe set and projection operators. We showcase our approach in two different applications inspired by rocket landing within a given landing window and safe autonomous blimp navigation, as well as on a physical robot testbed involving a wheeled robot navigating through a cluttered environment.
@@ II Problem Statement
Consider an autonomous system with state $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ that evolves according to dynamics
$\dot{\mathbf{x}}=f(\mathbf{x},u,d)$ (1)
where $u\in\mathcal{U}$ and $d\in\mathcal{D}$ are the control and disturbance of the system, respectively. $d$ can represent potential model uncertainties or an actual, adversarial exogenous input to the system. We assume the dynamics are uniformly continuous in $u$ and $d$, bounded, and Lipschitz continuous in $\mathbf{x}$ for fixed $u$ and $d$. Finally, let $\xi_{\mathbf{x},t}^{\mathbf{u},\mathbf{d}}(\tau)$ denote the system state at time $\tau$, starting from the state $\mathbf{x}$ at time $t$ under control signal $\mathbf{u}(\cdot)$ and disturbance signal $\mathbf{d}(\cdot)$ while following the dynamics (1). A control signal $\mathbf{u}(\cdot)$ is defined as a measurable function mapping from the time horizon to the set of admissible controls $\mathcal{U}$, and a disturbance signal is similarly defined. We additionally assume that the control and disturbance signals $\mathbf{u}(\cdot)$ and $\mathbf{d}(\cdot)$ are piecewise continuous in $t$. This assumption ensures that the system trajectory $\xi_{\mathbf{x},t}^{\mathbf{u},\mathbf{d}}$ exists and is unique and continuous for all initial states [11, 8].
In this work, we are interested in synthesizing potentially time-varying controllers $\pi:\left[0,T\right]\times\mathcal{X}\rightarrow\mathcal{U}$ that steer the system to reach (liveness) or avoid (safety) a given target set $\mathcal{L}\subseteq\mathcal{X}$, within the time horizon $\left[0,T\right]$. $\mathcal{L}$ can represent the goal region in the case of liveness or an unsafe region of the state space in the case of safety, e.g., obstacles for a navigation robot. Furthermore, we would like our controllers to consider other performance objectives while guaranteeing liveness/safety. We assume the performance objectives are encoded and optimized by an user-defined nominal controller $\pi_{\text{nom}}$. However, $\pi_{\text{nom}}$ may not necessarily ensure liveness (or safety). Our goal is to design $\pi$ to follow $\pi_{\text{nom}}$ to the extent possible, while guaranteeing liveness/safety requirement.
@@
Symbol Definition
$n\in\mathbb{Z}_{+}$ Number of states
$\mathcal{X}\subseteq\mathbb{R}^{n}$ State space
$\mathbf{x}\in\mathcal{X}$ System state
$\dot{\mathbf{x}}$ Time derivative of state $\mathbf{x}$
$\mathcal{U}$ Control space
$u\in\mathcal{U}$ Control input
$\mathcal{D}$ Disturbance space
$d\in\mathcal{D}$ Disturbance input
$f$ System dynamics
$\mathbf{u}(\cdot)$ Control signal (over time)
$\mathbf{d}(\cdot)$ Disturbance signal (over time)
$\xi_{\mathbf{x},t}^{\mathbf{u},\mathbf{d}}(\tau)$ State at time $\tau$ starting from state $\mathbf{x}$ at time $t$ under control and disturbance signals $\mathbf{u}(\cdot)$ and $\mathbf{d}(\cdot)$
$\mathcal{L}\subseteq\mathcal{X}$ Target set. Could represent goal or failure states.
$l:\mathcal{X}\!\rightarrow\!\mathbb{R}$ Target function. Signed distance to the target set.
$T$ Time horizon.
$\pi_{\text{nom}}$ Nominal controller
$V(\mathbf{x},t)$ Value function at state $\mathbf{x}$ and time $t$
$D_{t}V(\mathbf{x},t)$ Time derivative of the value function
$\nabla V(\mathbf{x},t)$ Spatial derivative of the value function
$\mathcal{G}_{\text{live}}(t)$ BRT of the target set in the liveness problem.
$\pi_{\text{live}}^{*}(\mathbf{x},t)$ Default liveness controller from the reachability analysis
$\mathcal{U}_{\text{live}}(\mathbf{x},t)$ Set of liveness preserving controls at state $\mathbf{x}$ and time $t$.
$h$ Projection operator for filtering
$\widetilde{\mathcal{U}}$ Projection set for filtering
$\gamma$ Constant used in the smooth blending filter.
$\mathcal{G}_{\text{unsafe}}$ BRT of the target set in the safety problem.
$\pi_{\text{safe}}^{*}(\mathbf{x})$ Default safety controller from the reachability analysis.
$\mathcal{U}_{\text{safe}}(\mathbf{x})$ Set of safety preserving controls.
TABLE I: A summary of all the symbols used in the paper.
@@
$f_{\text{rocket}}(\mathbf{x},u)=\dfrac{d}{dt}\begin{bmatrix}y\\ z\\ \theta\\ \dot{y}\\ \dot{z}\\ \dot{\theta}\end{bmatrix}=\begin{bmatrix}\dot{y}\\ \dot{z}\\ \dot{\theta}\\ \cos(\theta)u_{y}-\sin(\theta)u_{z}\\ \sin(\theta)u_{y}+\cos(\theta)u_{z}-g\\ \alpha u_{1}\end{bmatrix}$ (2)
@@
Here $y$, $z$ denotes the horizontal and vertical positions of the center of mass of the rocket, $\theta$ denotes the rocket’s heading, $u_{y}$ and $u_{z}$ denote the thrust in the $y$ and $z$ direction, respectively, and $g$ is the acceleration due to the gravity of Earth. The task of the liveness controller is to generate thrust inputs $u=[|u_{y}|<250,|u_{z}|<250]$ such that the system can reach a landing pad (target set) within one second. The target set is defined as the rectangular area =$\mathcal{L}=\{(y,z):|y|<20,0<z<20\}$. The corresponding target function is given by $l(\mathbf{x})=\max\{|y|-20,z-20,-z\}$, which approximates the signed distance to the target set. For this example, we assume no external disturbances are acting on the system.
@@
We first synthesize a nominal controller that attempts to take the system to $\mathcal{L}$. The nominal controller is a sampling-based model predictive controller (MPC), whose objective is to minimize the distance between center of mass position and the boundary of $\mathcal{L}$, and the control energy over the entire trajectory. Intuitively, the nominal controller can be considered as a performance control that tends to drive the system state to $\mathcal{L}$ while optimizing the amount of control effort. The MPC cost to be minimized over the entire trajectory is given as,
$\min_{u}$ $\sum_{i=1}^{t}||\left[u_{y_{i}},u_{z_{i}}\right]||_{2}+\sqrt{(|y_{i}|-20)^{2}+(|z_{i}|-20)^{2}}$ (3)
Here, the first term penalizes the control cost while the second term penalizes the distance from the target set. $i$ is the discrete timestep and ||.$||.||_{2}$ is the $L_{2}$ norm. The MPC controller is implemented in a receding horizon fashion with a horizon of 1000 steps. We use a first-order Euler discretization of the continuous dynamics in (2) with timestep, $\delta$ = 0.0005s. In our simulation studies, we will show that, in this case, such a nominal controller is insufficient to assure the liveness of our system. Liveness might be achievable using MPC, but designing such a controller would require more care and time. In this work, we will use a liveness filter to refine this nominal controller and produce a controller with the desired liveness properties.
@@ III-A Liveness and Safety Problems in HJ Reachability
In HJ reachability analysis, the liveness and safety problems are posed as optimal control problems; informally, both problems intend to find a control signal $\mathbf{u}(\cdot)$ that steers the system as “deep” into or as “far away” from the target set as possible. To capture this semantic, the minimum distance between the system and the target set $\mathcal{L}$ over the time horizon is defined as the objective of the optimal control problems. Let $l:\mathcal{X}\rightarrow\mathbb{R}$ be some bounded and Lipschitz continuous function whose sub-zero level is given by the target set: =$\mathcal{L}=\{\mathbf{x}:l(\mathbf{x})\leq 0\}$. Here, we present the reachability analysis for the liveness case and then comment on the safety case. Given $l$, the liveness problem is defined in (4):
$\begin{split}\max_{\mathbf{d}(\cdot)}\min_{\mathbf{u}(\cdot)}&\min_{\tau\in[t,T]}l(\xi_{\mathbf{x},t}^{\mathbf{u},\mathbf{d}}(\tau))\\ s.t.&\quad\dot{\mathbf{x}}=f(\mathbf{x},u,d)\end{split}$ (4)
where the minimum cost over time for the trajectory is defined as the cost function $J(\mathbf{x},t,\mathbf{u}(\cdot),\mathbf{d}(\cdot))=\min_{\tau\in[t,T]}l(\xi_{\mathbf{x},t}^{\mathbf{u},\mathbf{d}}(\tau))$. The above optimization problem (4) finds the minimum distance to the target set over the system trajectory, under the optimal control and the worst case disturbance. Thus, the system can reach the target set if and only if the minimum distance is less than 0, and this minimum distance is captured by the value function:
$V(\mathbf{x},t)=\max_{\mathbf{d}(\cdot)}\min_{\mathbf{u}(\cdot)}J(\mathbf{x},t,\mathbf{u}(\cdot),\mathbf{d}(\cdot))$ (5)
The value function (5) can be computed using dynamic programming, and it satisfies the following final value Hamilton-Jacobi-Isaacs Variational Inequality (HJI-VI) [20, 17]:
$\min\{D_{t}V(\mathbf{x},t)$ $+H(\mathbf{x},t,\nabla V(\mathbf{x},t)),l(\mathbf{x})-V(\mathbf{x},t)\}=0$ (6)
$\text{for}\ t\in\left[0,T\right]\;\text{and}\;V(\mathbf{x},T)=l(\mathbf{x})\quad$
Here, $D_{t}V(\mathbf{x},t)$ and $\nabla V(\mathbf{x},t)$ denote the temporal derivative and the spatial gradients of the value function $V(\mathbf{x},t)$, respectively. The Hamiltonian encodes how the control and disturbance interact with the system dynamics and is given as:
$H(\mathbf{x},t,\nabla V(\mathbf{x},t))=\min_{u\in\mathcal{U}}\max_{d\in\mathcal{D}}\nabla V(\mathbf{x},t)\cdot f(\mathbf{x},u,d)$ (7)
@@
$\mathcal{G}_{\text{live}}(t)\coloneqq\{\mathbf{x}:V(\mathbf{x},t)\leq 0\}$ (8)